Factor the trinomial
w^3-7w^2-18w

Answer by stanbon(75887)

Factor the trinomial
w^3-7w^2-18w
---
Common factor: w
---
= w(w^2 - 7w - 18)
---
= w(w^2-9w+2w-18)
---
= w(w(w-9)+2(w-9))
---
= w(w-9)(w+2)
================
Cheers,
Stan H.
================